# Lattice points on determinant surfaces and the spectrum of the automorphic Laplacian

### v1

@@ 12.3. Removal of the condition $nl\leq m$
Our next step is to capture the condition $nl\leq m$ by Lemma 2.4. So, we write
@@
$E_{1}:=\frac{1}{T}\mathop{\sum_{m}\sum_{n}}_{m\neq nl}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\left(\frac{m}{nl}\right)^{\sigma_{1}}\frac{1}{|\log(m/nl)|}\tilde{T}(m,n,C),$
$E_{2}:=\mathop{\sum_{m}\sum_{n}}_{m=nl}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\int\limits_{\sigma_{1}-iT}^{\sigma_{1}+iT}\tilde{T}(m,n,C)\frac{\textnormal{d}|s_{1}|}{|s_{1}|},$
@@
$E_{3}:=\frac{1}{T}\mathop{\sum_{m}\sum_{n}}_{m=nl}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\tilde{T}(m,n,C).$
@@
$E_{1}$ $\ll r_{1}^{\theta}X^{3/4}\frac{X^{1+\varepsilon}}{TH}MN,$ (12.17)
$E_{2}$ $\ll{r_{1}}^{\theta}X^{3/4+\varepsilon}\sqrt{MN},$
$E_{3}$ $\ll r_{1}^{\theta}\frac{X^{3/4+\varepsilon}}{T}\sqrt{MN}.$
@@
which yields the bound for $E_{1}$ once we recall Prop. 11.3.
@@
$U(M,N,C):=\sum_{\begin{subarray}{c}K\text{ dyadic}\\ \frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq K\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}\end{subarray}}\sum_{\kappa_{j}\sim K}\frac{\kappa_{j}^{2\sigma_{2}}}{\kappa_{j}^{2}}\Bigg{|}\sum_{\nu=0}^{\infty}\mathcal{V}(\kappa_{j},\nu,r_{1})\mathop{\sum_{m}\sum_{n}}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)$ (12.21)
×$\times\left(\frac{2\kappa_{j}}{\sqrt{\pi}X^{\frac{1}{2}}(mn)^{\frac{1}{4}}}\right)^{4\nu}\left(\frac{m}{nl}\right)^{s_{1}}(mn)^{-s_{2}}\rho_{j}(nr_{1})\rho_{j}(m)(mn)^{-i\kappa_{j}}\Bigg{|}$
@@
×$\times\sum_{\kappa_{j}\sim K}e^{-\pi\kappa_{j}}\left|\sum_{m}\sum_{n}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\left(\frac{m}{nl}\right)^{s_{1}}(mn)^{-s_{2}-\nu}\rho_{j}(nr_{1})\rho_{j}(m)(mn)^{-i\kappa_{j}}\right|$
@@
Here we have used Hecke multiplicativity (see (2.7)) to separate the factor $r_{1}$ in $\rho_{j}(nr_{1})$ in the manner of the proof of Lemma 11.2. Note that since $\sqrt{MNr_{1}}/C\ll 1$ (see (10.2)), the sum over $\nu$ is absolutely bounded. Now we execute the integrals in (LABEL:cs) over $s_{1}$ and $s_{2}$. Because of the decay of $\tilde{V}$, the integral over $s_{2}$ presents no difficulty and the integral over $s_{1}$ contributes a factor which is $O(\log X)$ if we choose $T=X^{10}$ (see (12.16)). Hence, by Prop. 12.2 with this choice of $T$, we have the bound
@@
If we did not use the large sieve (Lemma 2.16) here, we would have obtain the bound $U(M,N,C)\ll(MN)^{3/4},$ and putting this in (LABEL:cs), Prop. 12.2 shows that the error term can be no better than $O(X^{8/5+\varepsilon})$. The large sieve improves the bound to $O((MN)^{-1/4}(M+N))$ and this gives us a saving which is of size $(X/H)^{3/4}\asymp X^{3/8}$ for our final choice $H=\sqrt{X}$.
@@ 12.5. Estimation of $\mathcal{D}_{-,k}(M,N,C)$ for $k>0$
$||p_{k}||_{\infty}\ll\frac{1}{\eta}||p_{0}||_{\infty},\quad\text{and}\quad||p_{0}||_{\infty}\ll||A||_{\infty}\ll\frac{\left(\frac{\sqrt{MN}}{C}\right)^{2\nu}}{\left(\frac{\sqrt{MN}}{C}\right)^{1/2}}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}},$ (12.24)
@@
Combining (12.25), (12.8), (11.2), (10.2), ignoring the condition $nl\leq m$ by positivity and by Prop. 11.3, we have for any $k\geq 1$,
@@ 12.6. Estimation of $\mathcal{D}_{-,Z}(M,N,C)$
Bringing together (12.7), (12.23), (12.26), (12.27), and taking $\delta=1/20$ and $A=2$, we complete the proof of Prop. 12.1.
@@ 13.2. Decomposition of $\mathcal{D}_{-}(M,N,C)$
$\check{f}_{m,n,C,-,\mu}(\eta):=\frac{4}{\sinh(2\pi\eta)}\sum_{\nu=0}^{\infty}\frac{2^{-2\nu}}{\Gamma(1+2i\eta)\nu!(1+2i\eta)\dots(\nu+2i\eta)}I_{-,\mu}(m,n,C,\nu,\eta,r_{1}),$ (13.7)
@@
$\mathfrak{D}_{-,\mu}(M,N,C)$ $:=\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{\kappa_{j}\ll\frac{X^{1+\varepsilon}}{H}}\rho_{j}(nr_{1})\rho_{j}(m)\check{f}_{m,n,C,-,\mu}(\kappa_{j}).$ (13.8)
@@ 14.1. Decomposition of $\mathfrak{D}_{-,1}(M,N,C)$
Let us consider (i) first and we only estimate $\mathfrak{D}_{-,1,2}(M,N,C)$ below as the other one can be estimated in the same way. The proof is similar to the proof of Prop. 7.3 and we only give a sketch. First note that since $\eta\ll X^{1+\varepsilon}/H$, $u_{0}=\frac{2\eta}{\sqrt{\pi}X^{\frac{1}{2}}(mn)^{\frac{1}{4}}}\ll\frac{X^{\frac{1}{2}+\varepsilon}}{H}.$ We write the range $\left[\beta_{1}+\delta,\frac{X^{\frac{1}{2}+\varepsilon}}{H}\right]$ of $u_{0}$ as a union of dyadic subintervals of the form $[\beta_{1}+Y,\beta_{1}+2Y]$, where the dyadic variable $Y$ can take at most $O(\log X)$ values. Let us consider two cases separately: $\delta\leq Y\leq\frac{\beta_{1}}{2}$ and $Y>\frac{\beta_{1}}{2}$.
@@
by (14.7). We we apply Prop. 11.3 to estimate this. We first bring the $Y$-sum and the $\kappa_{j}$-sum outside of the sums over $m$ and $n$. Since $\beta_{1}$ depends on $m,n$, we extend the range of $Y$ and $\kappa_{j}$ by positivity, as follows. We have,
$Y>\frac{\beta_{1}}{2}\implies Y\geq\beta_{0},\quad\text{where}\quad\beta_{0}:=\frac{1}{1000}\frac{\sqrt{M}(MN)^{\frac{1}{4}}}{\sqrt{XN}}.$ (14.9)
@@
Now we consider part (ii). Here, $\eta$ is highly localized and Lemma 2.5 with $k=2$ suffices and we apply Prop. 13.3 to handle the sum over $\kappa_{j}$. We omit the details. ∎
@@ 14.3. Estimation of $\mathfrak{D}_{-,1,5}(M,N,C)$ and the weighted stationary phase analysis
$R_{1}:=\frac{|A(\beta_{1})|}{|h_{-}^{\prime}(\beta_{1})|}+\frac{|A(\alpha_{1})|}{|h_{-}^{\prime}(\alpha_{1})|},\text{ }R_{2}:=\frac{Q^{4}T}{P^{2}}\left(1+\frac{Q}{V}\right)^{2}\left(\frac{1}{(x_{0}-\alpha_{1})^{3}}+\frac{1}{(\beta_{1}-x_{0})^{3}}\right),\text{ and }R_{3}:=\frac{QT}{P^{\frac{3}{2}}}\left(1+\frac{Q}{V}\right)^{2},$ (14.11)
@@
×$\times\mathcal{V}(\kappa_{j},\nu,r_{1})\Big{[}\mathcal{L}(m,n,\kappa_{j},\nu,C)+O(R_{1}+R_{2}+R_{3})\Big{]}.$
@@ 14.3.1. Contribution of $G(M,N,C)$
by (12.2) and (14.1). This sum is similar to the sum in (12.13) except for the condition (14.13). In (12.13), the corresponding condition on $\kappa_{j}$ was automatic from the support of the function $V$. Here also the presence of the function $V$ in $\mathcal{L}(m,n,\kappa_{j},\nu,C)$ limits $\kappa_{j}$ to lie in the interval given by (12.11) and hence the above upper bound for $\kappa_{j}$ is automatically satisfied. So we have only one extra condition $\kappa_{j}>\pi m$ from which we need to separate the variables $m$ and $\kappa_{j}$. By Lemma 2.4,
@@
$X^{3/4}\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{\begin{subarray}{c}\kappa_{j}\\ \kappa_{j}=\pi m\end{subarray}}|\rho_{j}(nr_{1})\rho_{j}(m)|\frac{e^{-\pi\kappa_{j}}}{\kappa_{j}^{5/2}}$ (14.16)
@@ 14.3.2. Contribution of $H(M,N,C)$
since from the support of $V$, $u_{0}\asymp\frac{(MN)^{1/4}}{\sqrt{C}}$. Therefore, using the same analysis as in the proof of part (ii) of Prop. 14.2 and using (12.14), we conclude that the contribution of these two terms is
@@ 14.3.3. Contribution of $J(M,N,C)$
Thus we complete the proof of Prop. 14.3 by clubbing together (LABEL:illish), (LABEL:G-leading), (LABEL:kappa=m), (LABEL:error1), (14.19), (14.20), (14.17) and recalling (9.5). Combining (14.3), Prop. 14.2, and Prop. 14.3, Prop. 14.1 follows.
@@ 15. Esitimation of $\mathcal{C}_{-}(M,N,C)$ for $C\geq C_{0}$
Thus, proceeding as in the proof of Prop. 11.5 and by (10.4), we have,
@@
Bringing together (10.6) and the estimates for $\mathcal{D}_{+}(M,N,C)$, $\mathcal{C}_{+}(M,N,C)$, $\mathcal{D}_{-}(M,N,C)$ and $\mathcal{C}_{-}(M,N,C)$, i.e., Prop. 11.4, Prop. 11.6, Prop. 12.1, Prop. 13.1 and Prop. 15.1, we have proved Prop. 9.1.
@@
The proofs of the analogues of Prop. 11.4, Prop. 11.6, Prop. 12.1, Prop. 13.1, and Prop. 15.1 are along the same lines but with some minor changes. For example, the bound for $\mathcal{D}_{-}(M,N,C)$ in Prop. 12.1 (for $C\leq C_{1}$) is now replaced by $\left(r^{\theta}\left(X^{\frac{1}{4}+\varepsilon}\left(X/H\right)^{2}+X^{\frac{1}{3}}\left(X/H\right)^{\frac{3}{2}}\right)\right)$. To see this, note that much of the proof of the proposition carries over but now $\mathcal{L}(m,n,\kappa_{j},\nu,C)$ (see (LABEL:lt)) is changed to $\mathcal{L}^{\star}(m,n,\kappa_{j},\nu,C)$ where the factor $\left(\frac{{\pi}^{2}X}{{\eta}^{2}}\right)^{\frac{3}{4}}$ is now replaced by $\left(\frac{{\pi}^{2}X}{{\eta}^{2}}\right)^{\frac{1}{4}}$. Note that the main contribution to the bound in Prop. 12.1 comes from the term $E_{2}$ (see Prop. 12.2), $\mathcal{D}_{-,0}(M,N,C)$ (see (12.23)), and $\mathcal{D}_{-,k}(M,N,C)$ (see (12.26)). The term analogous to $E_{2}$ now contributes $O\left(r^{\theta}X^{\frac{1}{3}}\left(X/H\right)^{\frac{3}{2}}\right)$ and the same contribution also comes from the analogue of $\mathcal{D}_{-,0}(M,N,C)$. In both the cases, the exponent of $X$ changes from $\frac{3}{4}$ to $\frac{1}{4}$ and the exponent of $\kappa_{j}$ changes from $\frac{5}{2}$ to $\frac{3}{2}$. The analogue of $\mathcal{D}_{-,k}(M,N,C)$ now contributes $\left(r^{\theta}X^{\frac{1}{4}+\varepsilon}\left(X/H\right)^{2}\right)$ since the factor $\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}$ is now changed to $\left(\frac{C}{N^{2}l}\right)^{\frac{1}{4}}$. We obtain the bound $O\left(r^{\theta}X^{\varepsilon}\left(X^{{\frac{1}{4}}}\left(X/H\right)^{2}+X^{\frac{1}{4}}\left(X/H\right)^{\frac{3}{2}}\right)\right)$ for the analogue of $\mathcal{D}_{-}(M,N,C)$ in Prop. 13.1 (for $C>C_{1}$) as well. For the analogue of $\mathcal{D}_{+}(M,N,C)$ in Prop. 11.4, we now have the bound $O\left(r^{\theta}X^{\frac{1}{4}+\varepsilon}\left(X/H\right)^{2}\right)$. Together, the analogous of $\mathcal{C}_{-}(M,N,C)$ and $\mathcal{C}_{+}(M,N,C)$ in Prop.s 15.1 and 11.6 now contribute $O\left(X^{{\frac{1}{4}}+\varepsilon}{\left(X/H\right)}^{\frac{3}{2}}\right)$.
@@
Now, from Prop. 9.1, (9.1), and (9.2), we obtain the bound $O\left(r^{\theta}\frac{X^{2+\varepsilon}}{H}\right)$ for $R_{1,w}(X)$, and as discussed earlier, the bound for ${R^{\prime}}_{1,w}(X)$ is the same. Hence we obtain the bound $O\left(r^{\theta}X^{\varepsilon}\left(\frac{X^{2}}{H}+X\sqrt{H}\right)\right)$ for $T_{1,w}(X)$ (see (8.2)) and the same bound holds for $T_{2,w}(X)$. We have, $R_{3,w}(X)\ll r^{\theta}X^{\frac{1}{4}+\varepsilon}\left(X^{\frac{1}{4}}\left(X/H\right)^{2}+X^{\frac{1}{3}}\left(X/H\right)^{\frac{3}{2}}\right)$ by Prop. 9.2 and (8.4), and the same bound holds for ${R^{\prime}}_{3,w}(X)$. Hence, by (8.3), $T_{3,w}(X)\ll r^{\theta}X^{\frac{1}{4}+\varepsilon}\left(X^{\frac{1}{4}}\left(X/H\right)^{2}+X^{\frac{1}{3}}\left(X/H\right)^{\frac{3}{2}}+X\sqrt{H}\right)$ and the same bound holds for $T_{4,w}(X)$. Therefore, Prop. 3.4 follows from the above discussion and Prop. 5.6.

### v2

@@ 11.3. Removal of the condition $nl\leq m$
Our next step is to capture the condition $nl\leq m$ by Lemma the Perron formula (see [Dav80, Chap. 17]). So, we write
@@
$E_{1}=\frac{1}{T}\mathop{\sum_{m}\sum_{n}}_{m\neq nl}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\left(\frac{m}{nl}\right)^{\sigma_{1}}\frac{1}{|\log(m/nl)|}\tilde{T}(m,n,C),$
$E_{2}=\mathop{\sum_{m}\sum_{n}}_{m=nl}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\int\limits_{\sigma_{1}-iT}^{\sigma_{1}+iT}\tilde{T}(m,n,C)\frac{\textnormal{d}|s_{1}|}{|s_{1}|},$
@@
$E_{3}=\frac{1}{T}\mathop{\sum_{m}\sum_{n}}_{m=nl}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\tilde{T}(m,n,C).$
@@
$E_{1}\ll r_{1}^{\theta}X^{3/4}\frac{X^{1+\varepsilon}}{TH}MN,\quad E_{2}\ll{r_{1}}^{\theta}X^{3/4+\varepsilon}\sqrt{MN},\quad E_{3}\ll r_{1}^{\theta}\frac{X^{3/4+\varepsilon}}{T}\sqrt{MN}.$ (11.16)
@@
which yields the bound for $E_{1}$ once we recall Prop. 10.3.
@@
$U(M,N,C)=\sum_{\begin{subarray}{c}K\text{ dyadic}\\ \frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq K\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}\end{subarray}}\sum_{\kappa_{j}\sim K}\frac{\kappa_{j}^{2\sigma_{2}}}{\kappa_{j}^{2}}\Bigg{|}\sum_{\nu=0}^{\infty}\mathcal{V}(\kappa_{j},\nu,r_{1})\mathop{\sum_{m}\sum_{n}}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)$ (11.20)
$\times\left(\frac{2\kappa_{j}}{\sqrt{\pi}X^{\frac{1}{2}}(mn)^{\frac{1}{4}}}\right)^{4\nu}\left(\frac{m}{nl}\right)^{s_{1}}(mn)^{-s_{2}}\rho_{j}(nr_{1})\rho_{j}(m)(mn)^{-i\kappa_{j}}\Bigg{|}$
@@
$\times\sum_{\kappa_{j}\sim K}e^{-\pi\kappa_{j}}\left|\sum_{m}\sum_{n}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\left(\frac{m}{nl}\right)^{s_{1}}(mn)^{-s_{2}-\nu}\rho_{j}(nr_{1})\rho_{j}(m)(mn)^{-i\kappa_{j}}\right|$
@@
Here we have used Hecke multiplicativity (see (2.1)) to separate the factor $r_{1}$ in $\rho_{j}(nr_{1})$ in the manner of the proof of Lemma 10.2. Note that since $\sqrt{MNr_{1}}/C\ll 1$ (see (9.2)), the sum over $\nu$ is absolutely bounded. Now we execute the integrals in (LABEL:cs) over $s_{1}$ and $s_{2}$. Because of the decay of $\tilde{V}$, the integral over $s_{2}$ presents no difficulty and the integral over $s_{1}$ contributes a factor which is $O(\log X)$ if we choose $T=X^{10}$ (see (11.15)). Hence, by Prop. 11.2 with this choice of $T$, we have the bound
@@
If we did not use the large sieve (Lemma 2.6) here, we would have obtain the bound $U(M,N,C)\ll(MN)^{3/4},$ and putting this in (LABEL:cs), Prop. 11.2 shows that the error term can be no better than $O(X^{8/5+\varepsilon})$. The large sieve improves the bound to $O((MN)^{-1/4}(M+N))$ and this gives us a saving which is of size $(X/H)^{3/4}\asymp X^{3/8}$ for our final choice $H=\sqrt{X}$.
@@ 11.5. Estimation of $\mathcal{D}_{-,k}(M,N,C)$ for $k>0$
[…]
@@
Combining (11.24), (11.8), (10.2), (9.2), ignoring the condition $nl\leq m$ by positivity and by Prop. 10.3, we have for any $k\geq 1$,
@@ 11.6. Estimation of $\mathcal{D}_{-,Z}(M,N,C)$
Bringing together (11.7), (11.22), (11.25), (11.26), and taking $\delta=1/20$ and $A=2$, we complete the proof of Prop. 11.1.
@@ 12.2. Decomposition of $\mathcal{D}_{-}(M,N,C)$
$\check{f}_{m,n,C,-,\mu}(\eta)=\frac{4}{\sinh(2\pi\eta)}\sum_{\nu=0}^{\infty}\frac{2^{-2\nu}}{\Gamma(1+2i\eta)\nu!(1+2i\eta)\dots(\nu+2i\eta)}I_{-,\mu}(m,n,C,\nu,\eta,r_{1}),$ (12.7)
@@
$\mathfrak{D}_{-,\mu}(M,N,C)$ $=\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{\kappa_{j}\ll\frac{X^{1+\varepsilon}}{H}}\rho_{j}(nr_{1})\rho_{j}(m)\check{f}_{m,n,C,-,\mu}(\kappa_{j}).$ (12.8)
@@ 13.1. Decomposition of $\mathfrak{D}_{-,1}(M,N,C)$
Let us consider (i) first and we only estimate $\mathfrak{D}_{-,1,2}(M,N,C)$ below as the other one can be estimated in the same way. The proof is similar to the proof of Prop. 7.2 and we only give a sketch. First note that since $\eta\ll X^{1+\varepsilon}/H$, $u_{0}=\frac{2\eta}{\sqrt{\pi}X^{\frac{1}{2}}(mn)^{\frac{1}{4}}}\ll\frac{X^{\frac{1}{2}+\varepsilon}}{H}.$ We write the range $\left[\beta_{1}+\delta,\frac{X^{\frac{1}{2}+\varepsilon}}{H}\right]$ of $u_{0}$ as a union of dyadic subintervals of the form $\beta_{1}+Y,\beta_{1}+2Y]$, where the dyadic variable $Y$ can take at most $O(\log X)$ values. Let us consider two cases separately: $\delta\leq Y\leq\frac{\beta_{1}}{2}$ and $Y>\frac{\beta_{1}}{2}$.
@@
by (13.7). We we apply Prop. 10.3 to estimate this. We first bring the $Y$-sum and the $\kappa_{j}$-sum outside of the sums over $m$ and $n$. Since $\beta_{1}$ depends on $m,n$, we extend the range of $Y$ and $\kappa_{j}$ by positivity, as follows. We have,
$Y>\frac{\beta_{1}}{2}\implies Y\geq\beta_{0},\quad\text{where}\quad\beta_{0}=\frac{1}{1000}\frac{\sqrt{M}(MN)^{\frac{1}{4}}}{\sqrt{XN}}.$ (13.9)
@@
Now we consider part (ii). Here, $\eta$ is highly localized and Lemma 2.8 with $k=2$ suffices and we apply Prop. 12.3 to handle the sum over $\kappa_{j}$. We omit the details. ∎
@@ 13.3. Estimation of $\mathfrak{D}_{-,1,5}(M,N,C)$ and the weighted stationary phase analysis
$R_{1}=\frac{|A(\beta_{1})|}{|h_{-}^{\prime}(\beta_{1})|}+\frac{|A(\alpha_{1})|}{|h_{-}^{\prime}(\alpha_{1})|},\text{ }R_{2}=\frac{Q^{4}T}{P^{2}}\left(1+\frac{Q}{V}\right)^{2}\left(\frac{1}{(x_{0}-\alpha_{1})^{3}}+\frac{1}{(\beta_{1}-x_{0})^{3}}\right),\text{ and }R_{3}=\frac{QT}{P^{\frac{3}{2}}}\left(1+\frac{Q}{V}\right)^{2},$ (13.11)
@@
$\times\mathcal{V}(\kappa_{j},\nu,r_{1})\Big{[}\mathcal{L}(m,n,\kappa_{j},\nu,C)+O(R_{1}+R_{2}+R_{3})\Big{]}.$
@@ 13.3.1. Contribution of $G(M,N,C)$
by (11.2) and (13.1). This sum is similar to the sum in (11.12) except for the condition (13.13). In (11.12), the corresponding condition on $\kappa_{j}$ was automatic from the support of the function $V$. Here also the presence of the function $V$ in $\mathcal{L}(m,n,\kappa_{j},\nu,C)$ limits $\kappa_{j}$ to lie in the interval given by (11.11) and hence the above upper bound for $\kappa_{j}$ is automatically satisfied. So we have only one extra condition $\kappa_{j}>\pi m$ from which we need to separate the variables $m$ and $\kappa_{j}$. By the Perron formula,
@@ 13.3.2. Contribution of $H(M,N,C)$
since from the support of $V$, $u_{0}\asymp\frac{(MN)^{1/4}}{\sqrt{C}}$. Therefore, using the same analysis as in the proof of part (ii) of Prop. 13.2 and using (11.13), we conclude that the contribution of these two terms is
@@ 13.3.3. Contribution of $J(M,N,C)$
Thus we complete the proof of Prop. 13.3 by clubbing together (LABEL:illish), (LABEL:G-leading), (LABEL:kappa=m), (LABEL:error1), (13.19), (13.20), (13.17) and recalling (8.4). Combining (13.3), Prop. 13.2, and Prop. 13.3, Prop. 13.1 follows.
@@ 14. Esitimation of $\mathcal{C}_{-}(M,N,C)$ for $C\geq C_{0}$
Thus, proceeding as in the proof of Prop. 10.5 and by (9.4), we have,
@@
Bringing together (9.6) and the estimates for $\mathcal{D}_{+}(M,N,C)$, $\mathcal{C}_{+}(M,N,C)$, $\mathcal{D}_{-}(M,N,C)$ and $\mathcal{C}_{-}(M,N,C)$, i.e., Prop. 10.4, Prop. 10.6, Prop. 11.1, Prop. 12.1 and Prop. 14.1, we have proved Prop. 8.1.
@@
The proofs of the analogues of Prop. 10.4, Prop. 10.6, Prop. 11.1, Prop. 12.1, and Prop. 14.1 are along the same lines but with some minor changes. For example, the bound for $\mathcal{D}_{-}(M,N,C)$ in Prop. 11.1 (for $C\leq C_{1}$) is now replaced by $\left(r^{\theta}\left(X^{\frac{1}{4}+\varepsilon}\left(X/H\right)^{2}+X^{\frac{1}{3}}\left(X/H\right)^{\frac{3}{2}}\right)\right)$. To see this, note that much of the proof of the proposition carries over but now $\mathcal{L}(m,n,\kappa_{j},\nu,C)$ (see (LABEL:lt)) is changed to $\mathcal{L}^{\star}(m,n,\kappa_{j},\nu,C)$ where the factor $\left(\frac{{\pi}^{2}X}{{\eta}^{2}}\right)^{\frac{3}{4}}$ is now replaced by $\left(\frac{{\pi}^{2}X}{{\eta}^{2}}\right)^{\frac{1}{4}}$. Note that the main contribution to the bound in Prop. 11.1 comes from the term $E_{2}$ (see Prop. 11.2), $\mathcal{D}_{-,0}(M,N,C)$ (see (11.22)), and $\mathcal{D}_{-,k}(M,N,C)$ (see (11.25)). The term analogous to $E_{2}$ now contributes $O\left(r^{\theta}X^{\frac{1}{3}}\left(X/H\right)^{\frac{3}{2}}\right)$ and the same contribution also comes from the analogue of $\mathcal{D}_{-,0}(M,N,C)$. In both the cases, the exponent of $X$ changes from $\frac{3}{4}$ to $\frac{1}{4}$ and the exponent of $\kappa_{j}$ changes from $\frac{5}{2}$ to $\frac{3}{2}$. The analogue of $\mathcal{D}_{-,k}(M,N,C)$ now contributes $\left(r^{\theta}X^{\frac{1}{4}+\varepsilon}\left(X/H\right)^{2}\right)$ since the factor $\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}$ is now changed to $\left(\frac{C}{N^{2}l}\right)^{\frac{1}{4}}$. We obtain the bound $O\left(r^{\theta}X^{\varepsilon}\left(X^{{\frac{1}{4}}}\left(X/H\right)^{2}+X^{\frac{1}{4}}\left(X/H\right)^{\frac{3}{2}}\right)\right)$ for the analogue of $\mathcal{D}_{-}(M,N,C)$ in Prop. 12.1 (for $C>C_{1}$) as well. For the analogue of $\mathcal{D}_{+}(M,N,C)$ in Prop. 10.4, we now have the bound $O\left(r^{\theta}X^{\frac{1}{4}+\varepsilon}\left(X/H\right)^{2}\right)$. Together, the analogous of $\mathcal{C}_{-}(M,N,C)$ and $\mathcal{C}_{+}(M,N,C)$ in Prop.s 14.1 and 10.6 now contribute $O\left(X^{{\frac{1}{4}}+\varepsilon}{\left(X/H\right)}^{\frac{3}{2}}\right)$.
@@
Now, from Prop. 8.1 and (8.1), we obtain the bound $O\left(r^{\theta}\frac{X^{2+\varepsilon}}{H}\right)$ for $R_{1,w}(X)$, and as discussed earlier, the bound for ${R^{\prime}}_{1,w}(X)$ is the same. Hence we obtain the bound $O\left(r^{\theta}X^{\varepsilon}\left(\frac{X^{2}}{H}+X\sqrt{H}\right)\right)$ for $T_{1,w}(X)$ (see (7.22)) and the same bound holds for $T_{2,w}(X)$. We have, $R_{3,w}(X)\ll r^{\theta}X^{\frac{1}{4}+\varepsilon}\left(X^{\frac{1}{4}}\left(X/H\right)^{2}+X^{\frac{1}{3}}\left(X/H\right)^{\frac{3}{2}}\right)$ by Prop. 8.2 and (7.24), and the same bound holds for ${R^{\prime}}_{3,w}(X)$. Hence, by (7.23), $T_{3,w}(X)\ll r^{\theta}X^{\frac{1}{4}+\varepsilon}\left(X^{\frac{1}{4}}\left(X/H\right)^{2}+X^{\frac{1}{3}}\left(X/H\right)^{\frac{3}{2}}+X\sqrt{H}\right)$ and the same bound holds for $T_{4,w}(X)$. Therefore, Prop. 3.3 follows from the above discussion and Prop. 5.1.